# No more hard prompts: SoftSRV prompting for synthetic data generation

### v2

@@ 3.3 Empirical evaluation procedure
For the hard prompt baselines, we generate synthetic questions by querying the same backbone LLM using the relevant domain specific hard prompt template and questions taken from the benchmark training fold to populate the template. We conducted a search over temperature={1,2,4} and found a temperature of 2 to provide a balance of diversity and quality for these hard prompting methods.
For all methods, we generate 100K questions, repeating example questions from the training fold in a round-robin fashion. We then run a simple filtering, deduplication and subsampling pipeline to arrive at a target fine-tuning dataset size $N_{s}$. Details of this procedure are provided in Appendix A.1. We use $N_{s}$=50,000 for MBPP and GSM8K and $N_{s}$=20,000 for BoolQ.
3. Generate Answers. After generating the questions, all methods essentially follow the same procedure to generate answers using an off-the-self LLM. The only difference being, in the case of SoftSRV, we pass the question directly the the LLM without any domain specific prompting to preserve the domain agnostic nature. In the case of hard prompt baselines, we use a domain specific hard prompt template combined with the generated question to query the off-the-shelf LLM for an answer. Once we have full (questions, answer) fine-tuning examples generated, we run a decontamination process to remove any examples that may have been inadvertently leaked to the pretrained LLM, as is standard practice (details provided in Appendix A.2).
4. Fine-tune and Evaluate Downstream Model. Finally, for all methods, we use the generated (question, answer) pairs to fine-tune the target Gemma 2B model. We use a batch-size of 16 with sequence length 8192 and with a learning rate with linear warmup from 0 to 1e-6 over 100 steps, followed by a cosine annealing schedule. We evaluate the performance of these fine-tuned models on the test fold of the respective benchmark using the procedure and metric stated in Section 3.1.
3.4 Gemma 2B fine-tune & downstream evaluation
Figure 2: Full fine-tuning curves for the Gemma 2B model using different synthetically generated datasets as well as the non-synthetic training set.
Here, we present the performance of Gemma 2B fine-tuned on the generated synthetic datasets. Figures 2 plots the eval metrics for each dataset as a function of the number of fine-tuning steps times batch size normalized by training set size (essentially the training epoch modulo an additional constant factor due to sequence packing). These figures show that the model fine-tuned on data generated by the $\mathrm{SS}_{\mathrm{MC}}$ method generally outperforms the models fine-tuned on the data generated by the other methods. Comparing the hard prompt methods, the $\mathrm{HP}$ method outperforms the $\mathrm{HP}_{\mathrm{SR}}$ method on the GSM8K benchmark. For the MBPP benchmark, $\mathrm{HP}$ initially attains a similar performance as that of $\mathrm{HP}_{\mathrm{SR}}$, but both methods start to degrade as a function of fine-tuning steps. This may be due to a lack of diversity in the generated text given the small set of 384 training example questions for MBPP. We do not report the results of the $\mathrm{HP}_{\mathrm{SR}}$ method for BoolQ as it appears to struggle to produce reasonable outputs. The repeated self-critiques of $\mathrm{HP}_{\mathrm{SR}}$ on the lengthy input passages seems to lead it astray from the original intention of the task, producing questions asking for open-ended discussion of a passage rather than targeted true/false questions.
In our soft-prompting setting, two salient questions are what type of parametrization is effective for soft prompts and whether it is essential to have a contextual soft prompt that leverages the context token as opposed to a non-contextual soft prompt. Our empirical evaluations show that the non-contextual soft prompt method, $\mathrm{SS}_{\mathrm{NSP}}$, admits a lower performance than the other methods thereby indicating the effectiveness of contextual soft prompts. Then, by comparing the performance of the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{SS}_{\mathrm{MP2}}$ methods, we observe that the more expressive parametrization of $\mathrm{SS}_{\mathrm{MC}}$ is generally beneficial.
We also find that the model fine-tuned on the $\mathrm{SS}_{\mathrm{MC}}$ generated data outperforms the model fine-tuned on the non-synthetic training dataset for MBPP and GSM8K, indicating that, given enough of it, synthetic data can outperform even non-synthetic data. However, the same observation does not hold for BoolQ. The training set curve on BoolQ admits high variance, but it attains a higher accuracy overall. As discussed in Section 3.1, we expect generating questions for the BoolQ dataset to be more difficult stemming both from the fact that the data was generated by search queries thereby covering a wide range of topics, but also from containing much longer sequences extracted from the Wikipedia passages.
In Table 1, we show a comparison of the methods at a fixed epoch and at their max metric value attained over all tested fine-tuning steps. The fixed epoch was chosen to be the step where the model admits the largest fine-tuning performance on the training dataset. This clearly gives an advantage to the model fine-tuned on the training data, but it still provides a fair comparison between the synthetically generated datasets. We again find that the model performance when fine-tuned on the $\mathrm{SS}_{\mathrm{MC}}$ generated data outperforms the models fine-tuned on the other generated datasets at both the fixed epoch and at the max metric. For the the BoolQ dataset, the $\mathrm{HP}$ method admits a performance close to that of the $\mathrm{SS}_{\mathrm{MC}}$ method for the max metric, but its accuracy for the fixed epoch is considerably lower than that of $\mathrm{SS}_{\mathrm{MC}}$.
Table 1: Downstream task performance of Gemma 2B models fine-tuned on various sources of data. The number of non-synthetic examples used as seed data for the hard prompt and SoftSRV models is reported as $N_{r}$. The number of synthesized examples post-deduplication, $N_{s}$, is 50k apart from BoolQ where it is 20k. The base column reports the pre-trained model performance without fine-tuning. The ‘train’ column corresponds to the model fine-tuned on the non-synthetic examples.
task metric $N_{r}$ base train $\mathrm{HP}$ $\mathrm{HP}_{\mathrm{SR}}$ $\mathrm{SS}_{\mathrm{NSP}}$ $\mathrm{SS}_{\mathrm{MP2}}$ $\mathrm{SS}_{\mathrm{MC}}$
(fixed epoch)
MBPP pass@1 384 0.304 0.314 0.254 0.250 0.334 0.324 0.348
GSM8K accuracy 7,473 0.29 0.441 0.435 0.422 0.411 0.439 0.471
BoolQ accuracy 9,427 0.757 0.877 0.833 0.832 0.835 0.852
(max metric)
MBPP pass@1 384 0.304 0.314 0.326 0.326 0.338 0.348 0.378
GSM8K accuracy 7,473 0.29 0.441 0.456 0.435 0.424 0.439 0.478
BoolQ accuracy 9,427 0.757 0.877 0.851 0.838 0.845 0.854
3.5 Data scaling
Next, we analyze the effects of varying the number of synthetic examples generated. Specifically, we increase the number of generated examples from 20K, 50K, to 100K, testing both the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{HP}$ approaches for the BoolQ benchmark, given this benchmark appears the most challenging in terms of generating effective synthetic questions. Figure 3(a) shows that the performance of the model fine-tuned on $\mathrm{HP}$ stagnates between 20K and 50K synthetic examples, and only shows improvement when the number of generated examples reaches 100K. In contrast, the model fine-tuned on $\mathrm{SS}_{\mathrm{MC}}$ steadily improves as the generated dataset size increases. Going from 20K to 100K examples, $\mathrm{SS}_{\mathrm{MC}}$ performance increases at a 1.8 times faster rate relative to $\mathrm{HP}$ with respect to the max metric value. In particular, the $\mathrm{HP}$ method with 100K generated examples admits a comparable performance to that of the $\mathrm{SS}_{\mathrm{MC}}$ method with 50K generated examples.
Figure 3: In (a) we compare the BoolQ performance of Gemma 2B fine-tuned on data generated by $\mathrm{HP}$ and $\mathrm{SS}_{\mathrm{NSP}}$ as the number of generated examples increases. In (b) for the GSM8K benchmark, we show the Gemma 2B performance after fine-tuning on the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{HP}$ generated datasets against a sample of the same size from the curated TinyGSM dataset.
@@ 3.5 Data scaling
3.6 Comparison with curated hard prompt generated dataset (TinyGSM)
@@
Here, we compare against a high quality dataset, TinyGSM, within our fine-tuning evaluation framework. TinyGSM was expertly curated by Liu et al. (2023) for GSM8K-PAL, which is a program aided language model (PAL) variant of GSM8K that asks for questions to be answered in the form of Python functions. This has the advantage of enabling verification of the answer in a programmatic fashion. Liu et al. (2023) use GPT-3.5-turbo with hard prompts seeded with training questions from the original GSM8K dataset and from the GSM-IC dataset, which is a dataset crafted to incorporate irrelevant context in order to bolster model robustness (Shi et al., 2023). They use two types of prompts: the first asks to generate both questions and answers while the second requires two calls to the LLM to first generate a question and then an answer. Leveraging the fact that the solutions of the math word problems are written in Python, they then filter out any data that contains code that is not executable by a Python interpreter. They additionally filter out questions that do not contain numbers as this indicates flawed math problems.
@@
In order to evaluate the TinyGSM generated question in our setting, we randomly sample 100K questions from the publicly available TinyGSM dataset, then further subsample down to 50K using the same post-processing pipeline used by all other methods in our comparison. Finally, we generate answers, fine-tune and evaluate in the same fashion as the other hard prompt baselines. Figure 3(b) shows that a model fine-tuned on the $\mathrm{SS}_{\mathrm{MC}}$ dataset attains a performance close to that when fine-tuned on the sample from the $\mathrm{TinyGSM}$ dataset while the $\mathrm{HP}$ method lags behind both. TinyGSM performing closely to $\mathrm{SS}_{\mathrm{MC}}$ is encouraging given that the TinyGSM dataset is highly curated and tailored specifically for the GSM8K benchmark.
3.7 Distribution matching analysis
We next investigate the capacity of the SoftSRV and hard-prompting baseline to generate data samples that can match a target distribution of text. To assess the proximity between the generated and target distributions, we compute Mauve scores as in Pillutla et al. (2021), which can be viewed as a scalar summary of the divergence between the textual output of a generative model and a reference distribution. The mauve score is able to simultaneously measure both the model’s ability to avoid generating text outside the support of the target distribution (Type I error) and the ability to generate text with a large coverage of the target distribution support (Type II error). The method essentially computes a quantized distribution for the generated and target distribution and measures their KL divergence, producing a normalized score between 0 and 1, where 1 indicates the two distributions are maximally similar (for further details see Appendix A.3).
@@ 3.7 Distribution matching analysis
Table 2: The Mauve similarity scores of synthetic datasets computed with respect to the non-synthetic training and test fold of each dataset. The similarity computed between the train and test fold itself is also included in the final row of the table. Bolding indicates the max Mauve score achieved by a synthetic dataset.
@@
$\mathrm{HP}$ 0.622 0.933 0.663 0.463 0.914 0.784
$\mathrm{HP}_{\mathrm{SR}}$ 0.327 0.870 0.397 0.865
$\mathrm{SS}_{\mathrm{NSP}}$ 0.776 0.862 0.519 0.781 0.839 0.575
@@
In Table 2, we report the mauve scores for hard-prompt and SoftSRV-based synthetic datasets on the MBPP, GSM8K and BoolQ domains, measuring the distance to questions in both the train and test folds. In all cases, we find that a SoftSRV variant can synthesize text that is closer to the training and test dataset distribution than the hard-prompt approaches. This is perhaps unsurprising to see with respect to the train set since, by design, the SoftSRV prompts are trained to optimize the likelihood that the LLM sequentially decodes the text seen in these training examples, but is encouraging to see that the trend holds for the test set as well. Notably, among the SoftSRV variants, the $\mathrm{SS}_{\mathrm{MC}}$ method achieves the highest score on GSM8K and BoolQ, where it appears the additional flexibility afforded by its parameterization allows a very high-fidelity match to the target distribution. On the other hand, in the case of MBPP, which has relatively simple question distribution (see discussion in Section 3.1), the simplest SoftSRV parameterization ($\mathrm{SS}_{\mathrm{NSP}}$) attains the largest mauve score while the more complex SoftSRV variants produce lower similarity scores, perhaps due to the relatively small MBPP training set (only 384 examples). While the hard-prompt datasets generally have lower mauve scores than the SoftSRV counterparts, the $\mathrm{HP}$ variant is able to achieve high scores on GSM8K. Notably, in the case of the MBPP dataset, we measure very low similarity scores for $\mathrm{HP}_{\mathrm{SR}}$; we conjecture this may be due to the several rounds of rewriting, which results in straying further from the original seed question.
Finally, we note that the mauve score is not a direct indicator of downstream fine-tuning performance. However, it is another signal to provide further support for (parametrized) soft-prompting as a high-fidelity approach for text generation across domains.
4 Related Work
SoftSRV provides a novel contribution at the intersection of soft-prompting and synthetic text generation for targeted fine-tuning tasks. On each of these individual topics, there is a large and rapidly growing literature, which we touch on only briefly in the section.
4.1 Synthetic text for LLM training
As mentioned in the introduction, there is a significant recent body of work demonstrating the effectiveness of using a large language model to generate synthetic training data for a smaller model.
In terms for generating pre-training data, the collection of “Textbooks Are All You Need” white-papers outlines the process of training the Phi series of small LMs, using carefully curated prompt templates and seed data sources, and shows the large boost in quality that synthetic data can provide (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024). The Cosmopedia project (Ben Allal et al., 2024) conducts a similar study, while also providing detailed and transparent steps as well as open-sourcing the prompts and generated data. Li et al. (2024a) construct a pretraining set almost “from scratch” using a diverse set of hard prompts. To build these prompts, they assume access to taxonomies of fields/sub-fields/disciplines within an area of expertise and build a syllabus which culminates in a series of “key concepts”, each of which a large language model is then queried to generate a lesson on. Apart from focusing on pre-training rather than fine-tuning for a specific domain, these works require a non-trivial amount human effort for building and/or curating hard prompts for the generating LLM, which our effort seeks to minimize.
Finally, in the recent work of Lee et al. (2024), the authors propose an adaptive procedure where a large language model is used to generate targeted fine-tuning data for a small model based on examples that the small model has made mistakes on. The large model is prompted to rewrite variants of these questions using specialized per-domain hard prompts. Extending the general-purpose SoftSRV approach to an adaptive framework is a current area of investigation.
4.2 Soft-prompting & non-text data modalities
@@
As mentioned previously, the use of soft-prompting (or prompt-tuning) already has a significant history outside of targeted synthetic fine-tuning data generation. Primarily known for its use as a parameter efficient fine-tuning method (Lester et al., 2021), soft prompt training has also been used as a framework to learn or compress in-context instructions (Mu et al., 2023). Li et al. (2024b) has proposed training a secondary model to compress in-context instructions into soft prompt that is then prepended to the task hard prompt and is meant to generalize to even new tasks.
@@
While our study has been focused on generating synthetic text, there have been similar efforts in other modalities. For example, the ControlNet approach of Zhang et al. (2023) trains a diffusion model to produce images conditioned on contextual input, for example, image edges or 3D pose skeletons. Similarly, Gao et al. (2024) train a diffusion-based speech model to condition on a “simple speech representation” embedding to guide the generation of new synthetic speech data. Finally, in the case of text-to-image generators, there has been a significant amount of work in solving the “inverse” problem of mapping from an image back to a prompt (either hard prompt or a soft representation), so that one can more predictably generate synthetic images in certain styles (see Mahajan et al. (2024) and many references therein).
5 Conclusion
In this work, we have established the effectiveness of using contextual soft prompts, via the SoftSRV framework, for generating targeted synthetic fine-tuning data and its applicability across several different domains. We deploy the same SoftSRV pipeline across math, coding, and reasoning tasks, in each case we find SoftSRV is able to generate fine-tuning data that provide good downstream performance and fits well to the target distribution without any per-domain specialization needed. Given these results, there are several natural directions for further research. For example, we can view soft prompts as one particular class of parameter efficient tuning approach that is natural to leverage for data synthethesis, however, other approaches (such as LoRA) may be worth investigating as well. Finally, adapting the choice of context vector in order to generate the most effective synthetic data for improving the downstream model is a current and ongoing line of work.
Appendix A Empirical Evaluation Additional Details
A.1 Synthetic Question Postprocessing
For all methods (hard prompt and SoftSRV based), we query the model to generate 100K sequences. From these 100K, we first filter exact duplicates. From the remainder, we subsample to achieve the target size $N_{s}$. To encourage a diverse subsample, again for all methods, we cluster the data and select examples from each cluster randomly in round-robin fashion.
Concretely, using the scikit-learn library (Pedregosa et al., 2011), we apply MiniBatch $k$-means to vectorized data, which has been reduced in dimensionality using SVD. For all methods, we set the number of clusters for MiniBatch $k$-means to 700, reduced the dimensionality to 100 features and used sk.TfidfVectorizer for vectorization. Given the $k$-means clustering, we randomly select without replacement one point per cluster until $N_{s}$ questions are chosen.
A.2 Decontamination Process
A.3 Mauve Score Computation
We let $\mathcal{G}$ and $\mathcal{D}$ denote the generated and target/reference distributions, respectively, and compute the mauve scores for each synthetic dataset as follows. Using an embedding model, a vector representation is computed for each sequence in the synthetic and reference sets. These embeddings are then projected into a discrete set using $k$-means clustering, and a divergence curve is traced between the cluster distributions of $\mathcal{G}$ and $\mathcal{D}$, see Equation 1 in Pillutla et al. (2021). The mauve has value between 0 and 1 and corresponds to the area under the divergence curve, where higher scores are indicative of a closer match between $\mathcal{G}$ and $\mathcal{D}$. The same small LM serving as the context embedder for SoftSRV in Subsection 3.3 is used to compute per-token representations, which is then averaged to produce the sequence-level embedding. We use $k=32$ clusters for all domains as we found that $k=16$ or $k=64$ yields similar qualitative results.
A.4 Mixture of prompts with various values of $k$
Figure 4: Comparison of $\mathrm{SS}_{\mathrm{MPk}}$ variants for $k=1,2,4,8$ on MBPP. $\mathrm{SS}_{\mathrm{NSP}}$ corresponds to $k=1$.
Here, we conduct an initial exploration to measure the effect of changing the number of basis soft prompt matrices, $k$, of the $\mathrm{SS}_{\mathrm{MPk}}$ method. For the MBPP benchmark, Figure 4 shows the comparison in performance on MBPP for the $\mathrm{SS}_{\mathrm{MPk}}$ method for different values of $k$. We find that the model performance changes in a non-monotonic fashion as we increase the $k$ value and that $k=4$ attains the best model performance in the particular case of MBPP. While careful tuning of this parameter is necessary for optimal performance, it also represents a limitation of the method given a finite compute budget. In the main body we chose to report the results of the simplest $\mathrm{SS}_{\mathrm{MPk}}$, that is with $k=2$, uniformly for all datasets.
A.5 Diversification of the $\mathrm{HP}$ method
To arrive at the the $\mathrm{HP}$ method presented in the main body of the paper, we conducted multiple iterations of prompt engineering and template refinements. In particular, we found that asking the model to generate "10 different questions" per example question and using a higher decoding temperature was critical. We demonstrate this effect on the GSM8K benchmark in Figure 5 which compares the performance of a model fine-tuned on data generated by $\mathrm{HP}$ method to that of the model fine-tuned on its undiversified counterpart where the question template asks to generate one question per given example question and the default decoding temperature is used. We found similar results for the other datasets. Despite the relative small change in the settings, the difference in model performance is significant, demonstrating some of the idiosyncratic nature of hard-prompting approaches. We provide the template for the undiversified $\mathrm{HP}$ in Appendix B.
Figure 5: Performance of $\mathrm{HP}$ method with and without diversification.
Appendix B Hard Prompt Templates

### v3

@@ A.1 Benchmark Task Discussion
For the baselines, we cycle through questions found in the training fold and insert them into the relevant domain-specific prompt template and generate synthetic questions by querying the same backbone LLM used by SoftSRV. We conducted a search over temperature={1,2,4} and found a temperature of 2 to provide a balance of diversity and quality for the baseline approaches.
For all methods, we generate 100K questions, repeating example questions from the training fold in a round-robin fashion. We then run a simple filtering, deduplication and subsampling pipeline to arrive at a target fine-tuning dataset size $N_{s}$. Concretely, from these 100K, we first filter exact duplicates. Then, to encourage a diverse subsample, again for all methods, we cluster the data and select examples from each cluster randomly in round-robin fashion. That is, using the scikit-learn library (Pedregosa et al., 2011), we apply MiniBatch $k$-means to vectorized data, which has been reduced in dimensionality using SVD. For all methods, we set the number of clusters for MiniBatch $k$-means to 700, reduced the dimensionality to 100 features and used sk.TfidfVectorizer for vectorization. Given the $k$-means clustering, we randomly select without replacement one point per cluster until $N_{s}$ questions are chosen. We use $N_{s}$=50,000 for MBPP and GSM8K and $N_{s}$=20,000 for BoolQ.
A.3 Decontamination Process
A.4 Diversification of the $\mathrm{PT}$ method
Figure 5: Performance of $\mathrm{PT}$ method with and without diversification, i.e. the phrase generate "10 different questions".
To arrive at the the $\mathrm{PT}$ method presented in the main body of the paper, we conducted multiple iterations of prompt engineering and template refinements. In particular, we found that asking the model to generate "10 different questions" per example question and using a higher decoding temperature was critical. We demonstrate this effect on the GSM8K benchmark in Figure 5 which compares the performance of a model fine-tuned on data generated by $\mathrm{PT}$ method to that of the model fine-tuned on its undiversified counterpart where the question template asks to generate one question per given example question and the default decoding temperature is used. We found similar results for the other datasets. Despite the relative small change in the settings, the difference in model performance is significant, demonstrating some of the idiosyncratic nature of prompt template approaches. We provide the template for the undiversified $\mathrm{PT}$ in Appendix C.
@@ A.4 Diversification of the $\mathrm{PT}$ method
A.5 Comparison with Curated Prompt Template Generated Dataset (TinyGSM)
Figure 6: For the GSM8K benchmark, we show the Gemma 2B performance after fine-tuning on the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{PT}$ generated datasets against a sample of the same size from the curated TinyGSM dataset.
@@ A.5 Comparison with Curated Prompt Template Generated Dataset (TinyGSM)
Here, we compare against a high quality synthetic dataset, TinyGSM, within our fine-tuning evaluation framework. TinyGSM was expertly curated by (Liu et al., 2023) for GSM8K-PAL, which is a program aided language model (PAL) variant of GSM8K that asks for questions to be answered in the form of Python functions. This has the advantage of enabling verification of the answer in a programmatic fashion. (Liu et al., 2023) use GPT-3.5-turbo with prompt templates seeded with training questions from the original GSM8K dataset and from the GSM-IC dataset, which is a dataset crafted to incorporate irrelevant context in order to bolster model robustness (Shi et al., 2023). They use two types of prompts: the first asks to generate both questions and answers while the second requires two calls to the LLM to first generate a question and then an answer. Leveraging the fact that the solutions of the math word problems are written in Python, they then filter out any data that contains code that is not executable by a Python interpreter. They additionally filter out questions that do not contain numbers as this indicates flawed math problems.
@@
In order to evaluate the TinyGSM generated question in our setting, we randomly sample 100K questions from the publicly available TinyGSM dataset, then further subsample down to 50K using the same post-processing pipeline used by all other methods in our comparison. Finally, we generate answers, fine-tune and evaluate in the same fashion as the other prompt template baselines. Figure 6 shows that a model fine-tuned on the $\mathrm{SS}_{\mathrm{MC}}$ dataset attains a performance close to that when fine-tuned on the sample from the $\mathrm{TinyGSM}$ dataset while the $\mathrm{PT}$ method lags behind both. TinyGSM performing closely to $\mathrm{SS}_{\mathrm{MC}}$ is encouraging given that the TinyGSM dataset is highly curated and tailored specifically for the GSM8K benchmark.
A.6 Mixture of Prompts with Various Values of $k$
Figure 7: Comparison of $\mathrm{SS}_{\mathrm{MPk}}$ variants for different values of $k$. $\mathrm{SS}_{\mathrm{NP}}$ corresponds to $k=1$.
Here, we conduct a exploration to measure the effect of changing the number of basis embedding matrices, $k$, of the $\mathrm{SS}_{\mathrm{MPk}}$ method.
Figure 7 shows the comparison in performance on various benchmarks across different values of $k$. We see an inherent trade-off as we increase $k$, which increases the capacity of the contextual embedding but also then require training more parameters. For MBPP we see that the value $k=4$ appears to be optimal, while for GSM8K and BOOLQ we see performance peaks at $k=2$. We fix $k=2$ throughout the evaluations in the main paper as a good general choice across different tasks.
A.7 Mauve Score Computation
We let $\mathcal{G}$ and $\mathcal{D}$ denote the generated and target/reference distributions, respectively, and compute the mauve scores for each synthetic dataset as follows. Using an embedding model, a vector representation is computed for each sequence in the synthetic and reference sets. These embeddings are then projected into a discrete set using $k$-means clustering, and a divergence curve is traced between the cluster distributions of $\mathcal{G}$ and $\mathcal{D}$, see Equation 1 in (Pillutla et al., 2021). The mauve has value between 0 and 1 and corresponds to the area under the divergence curve, where higher scores are indicative of a closer match between $\mathcal{G}$ and $\mathcal{D}$. The same small LM serving as the context embedder $E$ for SoftSRV in Subsection 3.3 is used here to compute per-token representations, which is then averaged to produce the sequence-level embedding. We use $k=32$ clusters for all domains as we found that $k=16$ or $k=64$ yields similar qualitative results.
In Table 5 we additionally report the mauve score computed when using the training set as a reference sample.
@@ A.7 Mauve Score Computation
Table 5: The mauve similarity scores of synthetic datasets computed with respect to the non-synthetic training and test fold of each dataset. The similarity computed between the train and test fold itself is also included in the final row of the table.
@@
$\mathrm{PT}$ 0.622 0.933 0.663 0.463 0.914 0.784
$\mathrm{PT}_{\mathrm{SR}}$ 0.327 0.870 0.397 0.865
$\mathrm{SS}_{\mathrm{NP}}$ 0.776 0.862 0.519 0.781 0.839 0.575
@@
Appendix B Related Work on Prompt-tuning & Non-text Data Modalities
@@
Prompt-tuning is primarily known for its use as a parameter efficient fine-tuning method (Lester et al., 2021; Li & Liang, 2021). It has also been used as a framework to learn or compress in-context instructions (Mu et al., 2023). Recently, Li et al. (2024b) proposed training a secondary model to compress in-context instructions into soft prompt that is then prepended to the task prompt templates.
@@
While our study has been focused on generating synthetic text, there have been similar efforts in other modalities. For example, the ControlNet approach of Zhang et al. (2023) trains a diffusion model to produce images conditioned on contextual input, for example, image edges or 3D pose skeletons. Similarly, Gao et al. (2024) trains a diffusion-based speech model to condition on a “simple speech representation” embedding to guide the generation of new synthetic speech data. Finally, in the case of text-to-image generators, there has been a significant amount of work in solving the “inverse” problem of mapping from an image back to a prompt, so that one can more predictably generate synthetic images in certain styles (see (Mahajan et al., 2024) and many references therein).
Appendix C Natural Language Prompt Templates